OpenStudy (anonymous):

Solve formula for the indicated letter. Assume that all variables represent nonnegative numbers. A = A0(1 – r)2 for r (A business formula)

OpenStudy (anonymous):

A = A0(1 – r)^2 for r A^(1/2)= A0(1 – r) A^(1/2) /A0= 1 – r r=1-[A^(1/2) /A0] ans